OpenStudy (anonymous):

what's the derivative of the square root of 9-(x^3)?

OpenStudy (anonymous):

(-3/2)x^2 / sqr root of 9-x^3

OpenStudy (anonymous):

\[\frac{d}{dx}\sqrt{9-x^3}\]?

OpenStudy (anonymous):

is it 1/2 divided by sqr root 9-x^3

OpenStudy (anonymous):

\[\frac{1}{2\sqrt{9-x^3}}\times -3x^2=-\frac{3x^2}{2\sqrt{9-x^3}}\]